A small pump can fill a pool in one hour less than twice as long as a large pump. If the two pumps were working together, they would fill the pool in 72 minutes (6/5 of an hour). How long would it take the small pump working alone to fill the pool? How long would it take the large pump?
Answer by richwmiller(17219)

s=2b-1
1/s+1/b=5/6
1/(2b-1)+1/b=5/6
(3b-1)/(b(2b-1)) = 5/6
b=2
b=.3
check to see if either solution is extraneous